# PV-S3111Photovoltaic-Semi Supervised Segmentation : Advancing Automatic Photovoltaic Defect Detection using Semi-Supervised Semantic Segmentation of Electroluminescence Images

### v3

@@ Abstract
Photovoltaic (PV) systems allow us to tap into all abundant solar energy, however they require regular maintenance for high efficiency and to prevent degradation. Traditional manual health check, using Electroluminescence (EL) imaging, is expensive and logistically challenging which makes automated defect detection essential. Current automation approaches require extensive manual expert labeling, which is time-consuming, expensive, and prone to errors. We propose PV-S3 (Photovoltaic-Semi Supervised Segmentation), a Semi-Supervised Learning approach for semantic segmentation of defects in EL images that reduces reliance on extensive labeling. PV-S3 is a Deep learning model trained using a few labeled images along with numerous unlabeled images. We introduce a novel Semi Cross-Entropy loss function to deal with class imbalance. We evaluate PV-S3 on multiple datasets and demonstrate its effectiveness and adaptability. With merely 20% labeled samples, we achieve an absolute improvement of 9.7% in IoU, 13.5% in Precision, 29.15% in Recall, and 20.42% in F1-Score over prior state-of-the-art supervised method (which uses 100% labeled samples) on UCF-EL dataset (largest dataset available for semantic segmentation of EL images) showing improvement in performance while reducing the annotation costs by 80%. For more details, visit our GitHub repository: https://github.com/abj247/PV-S3.
Keywords: Photovoltaic Modules; Machine learning; Defect Detection; Semantic Segmentation; Semi-Supervised Learning
@@ 1 Introduction
To overcome this limitation, we propose PV-S3 (Photovoltaic-Semi Supervised Segmentation), a semi-supervised deep learning framework for semantic segmentation in PV module defect detection, which efficiently utilize both labeled and unlabeled EL images. PV-S3 reduces the reliance on extensive labeled data while addressing the scalability issues in large-scale solar installations. It is based on mean-teacher approach [42] and efficiently leverages unlabeled data. The proposed method enhances model accuracy and generalization by enforcing consistency between the predictions of a student model and a temporally averaged teacher model. Such an approach is crucial for reducing reliance on extensive labeled datasets, addressing a key challenge in the semantic segmentation of PV module defects. To the best of our knowledge there is no current existing work in Photovoltaics making use of semi-supervised learning approaches or reduced labeled data for defect detection.
@@
Figure 1: An overview of the proposed semi-supervised approach for PV defect detection: It utilizes both labeled and unlabelled images in a semi-supervised framework (SSL) using supervised and unsupervised consistency losses to train the defect detection model (top row) and during testing the trained model predicts the defect class for each pixel in EL images (bottom row).
@@
    We demoonstrate that the proposed approach outperforms existing state-of-the-art fully supervised methods with merely 20% annotations.
@@ 2 Related Work
Our research introduces a novel approach to PV module defect detection, leveraging semantic segmentation and a semi-supervised learning framework. Unlike prior works, our method efficiently incorporates both labeled and unlabeled data, enhancing model performance without heavy reliance on extensive labeled datasets.
@@ 3.2.2 SemiCE as Consistency Loss
$\mathcal{L}_{\text{pos}}=-\frac{1}{N_{\text{U}}}\sum_{i=1}^{N_{\text{U}}}\sum_{j=1}^{H\times W}\begin{cases}Y_{\text{L},ij}\cdot\log\left(\frac{\exp(P_{\text{$\theta$},ij})}{\sum_{k=1}^{H\times W}\exp(P_{\text{$\theta$},ik})}\right),&\text{if }Y_{\text{L},ij}=1\text{ and }P_{\text{$\theta$},ij}\geq t_{\text{pos}}\\ 0,&\text{otherwise}\end{cases}$ (2)
@@
$\mathcal{L}_{\text{neg}}=-\frac{1}{N_{\text{U}}}\sum_{i=1}^{N_{\text{U}}}\sum_{j=1}^{H\times W}\begin{cases}(1-Y_{\text{L},ij})\cdot\log\left(\frac{\exp(-P_{\text{$\theta$},ij})}{\sum_{k=1}^{H\times W}\exp(-P_{\text{$\theta$},ik})}\right),&\text{if }Y_{\text{L},ij}=0\text{ and }P_{\text{$\theta$},ij}<-t_{\text{neg}}\\ 0,&\text{otherwise}\end{cases}$ (3)
The final Semi Cross Entropy Loss $\mathcal{L}_{\text{SemiCE}}$ is the sum of $\mathcal{L}_{\text{pos}}$ and $L_{\text{neg}}$: The SemiCE loss is the same as the Consistency loss ($\mathcal{L}_{\text{cons}})$ for the student-teacher framework where this loss component tries to make the predictions of the student model consistent.
$\mathcal{L}_{\text{cons}}=\mathcal{L}_{\text{SemiCE}}=\mathcal{L}_{\text{pos}}+\mathcal{L}_{\text{neg}}$ (4)
In the equations, $N_{\text{U}}$ represents the total number of unlabeled images, $P_{\text{$\theta$}_{i}(j)}$ represent the logits predicted by the model, targets are the ground truth labels, $Y_{\text{L}_{i}(j)}$ is the ground truth label for the $i$-th sample and $j$-th class, and $H\times W$ represents the spatial dimensions of the image.
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sup}}+\lambda\mathcal{L}_{\text{cons}}$ (5)
where $\lambda$ is the weighting coefficient that determines the importance of the consistency loss relative to the supervised loss.
@@ 4 Experiments and Evaluation
Evaluation Metric
We evaluate the proposed method using Mean Intersection over Union (IoU), precision, recall, and F1 score. Mean IoU measures the overlap between the predicted segmentation masks and the ground truth annotations, providing an overall assessment of the segmentation accuracy. Precision represents the ratio of correctly predicted defect pixels to all predicted defect pixels, while recall measures the proportion of correctly predicted defect pixels out of all ground truth defect pixels. F1 score combines precision and recall, providing a balanced measure of the model’s overall performance. To clearly define these evaluation metrics, we provide their mathematical formulations:
  • 1.
    Precision is calculated as:
    Precision=True Positives (TP)True Positives (TP)+False Positives ($\text{Precision}=\frac{\text{True Positives (TP)}}{\text{True Positives (TP)}+\text{False Positives (FP)}}$ (6)
    where TP represents the number of correctly predicted positive pixels, and FP represents the number of pixels incorrectly predicted as positive.
  • 2.
    Recall is calculated as:
    Recall=True Positives (TP)True Positives (TP)+False Negatives ($\text{Recall}=\frac{\text{True Positives (TP)}}{\text{True Positives (TP)}+\text{False Negatives (FN)}}$ (7)
    where FN represents the number of positive pixels that were not correctly predicted.
  • 3.
    F1 Score is the harmonic mean of precision and recall:
@@
  • 4.
    Mean Intersection over Union (IoU) is calculated as:
    Mean $\text{Mean IoU}=\frac{1}{N_{\text{classes}}}\sum_{i=1}^{N_{\text{classes}}}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}$ (9)
    where $N_{\text{classes}}$ is the number of classes, and $\text{TP}_{i}$, $\text{FP}_{i}$, and $\text{FN}_{i}$ are the true positives, false positives, and false negatives for class $i$, respectively.
Here, true positives (TP) are pixels correctly identified as defects, false positives (FP) are pixels incorrectly identified as defects, and false negatives (FN) are defect pixels that were missed by the model. These equations provide a quantitative basis for evaluating and comparing the performance of our segmentation model.
@@ 4.1 Quantitative Evaluation
Regarding the performance on UCF-EL dataset with the semi-supervised learning scenario with 20% labeled data, we record average values for metrics as IoU of 67%, precision of 72.50%, recall of 88.90%, and F1-score of 78.17% whereas with 20% labeled data in the supervised learning yielded mean IoU of 8.76%, precision of 25.79%, recall of 26.27% and F1-score of 14.73%. Based on the results we can observe that semi-supervised learning with PV-S3 has yielded significant improvement in results as compared to fully supervised learning with same amount of labeled dataset. The scores with the baseline model of supervised learning with 20% labeled data and semi-supervised learning with 20% labeled data has been summarized in Table  1. To further evaluate the robustness of PV-S3, we perform three experimental runs with different subsets of labeled samples which are randomly selected from the training dataset. The mean and standard deviation of performance metrics for each defect class are shown in Table  2. We observe that the proposed SSL approach significantly outperform the supervised baseline in all metrics consistently, demonstrating its robustness.
IoU Precision Recall F1-Score
@@
IoU Precision Recall F1-Score
@@
With 100% labeled data using fully supervised learning, the Deeplabv3+ model achieves mean metrics of IoU: 70.84%, precision: 80.12%, recall: 86.29%, and F1-score: 82.03% (Table  3). Comparing these results with our 20% labeled semi-supervised learning approach, PV-S3 yields comparable results to fully supervised learning, with even higher precision. Leveraging 100% labeled data in the semi-supervised framework with PV-S3 significantly improves performance, yielding mean IoU: 74.75%, precision: 79.96%, recall: 91.30%, and F1-score: 84.24%. This improvement underscores the benefits of a larger dataset in better understanding and segmenting defects in PV module EL images.
IoU Precision Recall F1-Score
@@
Table 3: Comparison of PV-S3 with our fully supervised method on UCF-EL dataset: The table shows the performance of different defect classes under Semi Supervised Learning (SSL) with 100% labeled images, and Fully Supervised Learning (FSL) with 100% labeled images. The metrics evaluated include IoU, Precision, Recall, and F1-Score, facilitating a direct comparison of the impact of labeling extent on model performance.
Baselines Labeled Data (%) IoU (%) Precision (%) Recall (%) F1-Score (%)
@@
Comparisons with existing Fully Supervised learning on UCF-EL dataset, performed by [17], across key metrics are summarized in Table  4. We observe improved performance with merely 20% labeled samples and significant enhancement in IoU, Precision, and F1-Score using all labeled samples. Our approach, utilizing just 20% of labeled samples, yields a notable absolute enhancement, including a 9.7% increase in IoU, a 13.5% rise in Precision, a 29.15% boost in Recall, and a 20.42% improvement in F1-Score compared to the previous state-of-the-art supervised method, reducing annotation expenses by 80%.
Experiment IoU (%) Precision (%) Recall (%) F1-Score (%)
@@
On the Combined Solar Benchmark Dataset (CSB-Dataset) we evaluate PV-S3 on all three subsets with three major defect classes Crack, Gridline and Inactive. The results for four key metrics, IoU, Precision, Recall and F1-Score, are summarized in Table  5. We observe competitive performance with PV-S3 using only 20% labeled samples across all metrics when compared with fully supervised method using 100% labeled samples. We also compare the performance of PV-S3 with existing methods on these datasets. This comparison is shown in Table  6. The performance of PV-S3 is compared with various baseline models which include PSPNet, UNet-12 and UNet-25 using both 20% (Our-SSL) and 100% (Our-FSL) labels. We observe that conventional semantic segmentation models like PSPNet and UNet variants which are UNet-12 and UNet-25 performed poorly in terms of the Intersection of Union as compared to Deeplab variants. In terms of recall score, UNet-25 has yielded the highest recall score of 60.66% while PSPNet yielded the lowest score of 43.0%. We observe that PV-S3 outperforms other baseline models in terms of all metric scores. Using merely 20% of labeled images for training we observed an increase of 30% in IoU with a reduction in 80% annotation cost compared to the state-of-the-art model on CSB-Dataset, further using our model with 100% labeled images an increase of 35.67% in IoU and 0.33% increase in recall was observed with state-of-the-art model on CSB-Dataset.
Baseline Models Labeled Data (%) IoU (%) Precision (%) Recall (%) F1-Score (%)
@@ 4.2 Discussion and Analysis
IoU Precision Recall F1-Score
Defect Class MSE SemiCE-S1 SemiCE-S2 MSE SemiCE-S1 SemiCE-S2 MSE SemiCE-12 SemiCE-S2 MSE SemiCE-S1 SemiCE-S2
@@
Detecting interconnect defects presents a significant challenge due to class imbalance: these defects appear in a small proportion of images but constitute a substantial percentage of the pixel distribution. To effectively address this, we employ the Semi Cross-Entropy (SemiCE) loss function, which is specifically designed to handle class imbalances by adjusting the weighting of positive and negative samples during training. The SemiCE loss function incorporates adaptive thresholds for positive and negative predictions. By setting a lower positive threshold, we include more predictions with lower confidence as positive samples. This approach increases the recall of defect detection by identifying more potential defects but may introduce additional false positives. Conversely, a higher positive threshold requires greater confidence for a prediction to be considered positive, enhancing precision but potentially missing some defects. The negative threshold focuses on confidently negative predictions. Setting an appropriate negative threshold ensures that only predictions with high confidence of being negative contribute to the loss. This prevents the model from being influenced by uncertain negative samples, which could degrade performance. By carefully tuning these positive and negative thresholds, the SemiCE loss function balances the influence of classes during training. This adjustment allows the model to handle class imbalance effectively, improving its ability to detect interconnect defects with higher accuracy and robustness.
@@
In our experiments, we varied both the SemiCE positive and negative thresholds to observe their impact on Interconnect defect class performance. Performance with different thresholds is shown in Table 8, with original results in Table  1 using initial thresholds of 0.6 (positive) and 0 (negative), chosen for their balanced nature. Through experimentation, we aimed to identify threshold values optimizing the metrics while maintaining model consistency, shedding light on the intricate relationship between prediction confidence and model performance. The best threshold values achieved were 0.2 (positive) and 0.4 (negative), yielding improved performance for the Interconnect defect class with an IoU of 64%, precision of 74%, recall of 83%, and F1-Score of 78%.
Threshold Crack Contact Interconnect Corrosion
Pos, Neg IoU Prec Rec F1 IoU Prec Rec F1 IoU Prec Rec F1 IoU Prec Rec F1
0.0, 0.0 89 95 93 94 51 65 71 68 50 71 65 78 92 98 94 96
0.2, 0.4 89 96 92 94 53 63 78 70 64 74 83 78 81 91 88 89
0.3, 0.2 89 98 91 94 52 61 77 68 44 50 79 61 67 69 96 80
0.4, 0.2 91 97 93 95 52 64 73 68 45 48 81 62 76 77 98 86
0.5, 0.5 91 98 93 95 53 64 75 69 46 46 97 63 85 87 97 92
0.6, 0.0 89 98 90 94 49 57 79 66 40 43 89 58 90 92 97 95
@@
We evaluated PV-S3’s performance using different proportions of labeled data (5% to 20%) and compared it with fully supervised learning using 100% labeled data. Results in Figure  8 indicate that semi-supervised learning with 5% labeled data initially underperforms compared to fully supervised learning. However, performance notably improves at 10% labeled data and continues to increase gradually up to 20%, approaching scores achieved with 100% labeled data for IoU, Recall, and F1 score. Notably, recall scores for semi-supervised learning even exceed those of fully supervised learning. Further we observe that the recall score for the semi-supervised model trained with 20% labeled data and 80% unlabeled data exceeds that of the model trained with 100% labeled data. Usually, semi-supervised learning methods which utilized unlabelled samples outperform fully supervised methods [48, 39, 49], This counterintuitive result can be attributed to the benefits of semi-supervised learning in leveraging unlabeled data to improve model generalization. The inclusion of unlabeled data acts as a regularizer, preventing the model from overfitting to the labeled data, which might contain noise or biases. In contrast, training on 100% labeled data without the additional unlabeled data may lead to overfitting, especially if the labeled data is limited in diversity.By incorporating a large amount of unlabeled data, the semi-supervised model learns more robust feature representations, capturing the underlying structure of the data more effectively. The semi-supervised approach allows the model to generalize better to unseen data, resulting in higher recall by correctly identifying more true positive instances of defects.
@@
We analyze the training process of PV-S3 with epochs and compare with the supervised approach. The plots for different metrics are shown in Figure  9. Here we observe that for supervised learning with only 20% images (SL (20%)), due to less number of images and class imbalance, the model is not able to learn the features accurately and hence it has very low pixel accuracy and high loss values and have low values of precision and recall as well. For PV-S3, we observe that it reaches high values in pixel accuracy, loss, precision, and recall as training progresses.
Figure 9: Analyzing model training: The top row plots show the variation of pixel accuracy (left), loss (middle), and IoU (right) with varying training epochs, and the bottom row shows the performance metrics variation with epochs for supervised and semi-supervised learning with 20% images for precision (left), recall (middle), and F1-score (right).

### v4

# Advancing Automatic Photovoltaic Defect Detection using Semi-Supervised Semantic Segmentation of Electroluminescence Images

@@ Abstract
Photovoltaic (PV) systems allow us to tap into all abundant solar energy, however they require regular maintenance for high efficiency and to prevent degradation. Traditional manual health check, using Electroluminescence (EL) imaging, is expensive and logistically challenging which makes automated defect detection essential. Current automation approaches require extensive manual expert labeling, which is time-consuming, expensive, and prone to errors. We propose PV-S3 (Photovoltaic-Semi-supervised Semantic Segmentation), a Semi-Supervised Learning approach for semantic segmentation of defects in EL images that reduces reliance on extensive labeling. PV-S3 is an artificial intelligence (AI) model trained using a few labeled images along with numerous unlabeled images. We introduce a novel Semi Cross-Entropy loss function to deal with class imbalance. We evaluate PV-S3 on multiple datasets and demonstrate its effectiveness and adaptability. With merely 20% labeled samples, we achieve an absolute improvement of 9.7% in mean Intersection-over-Union (mIoU), 13.5% in Precision, 29.15% in Recall, and 20.42% in F1-Score over prior state-of-the-art supervised method (which uses 100% labeled samples) on University of Central Florida-Electroluminescence (UCF-EL) dataset (largest dataset available for semantic segmentation of EL images) showing improvement in performance while reducing the annotation costs by 80%. For more details, visit our GitHub repository: https://github.com/abj247/PV-S3.
Keywords: Photovoltaic Modules; Machine learning; Defect Detection; Semantic Segmentation; Semi-Supervised Learning; Artificial Intelligence Model
@@ 1 Introduction
To overcome this limitation, we propose PV-S3 (Photovoltaic-Semi Supervised Semantic Segmentation), a semi-supervised deep learning framework for semantic segmentation in PV module defect detection, which efficiently utilize both labeled and unlabeled EL images. PV-S3 reduces the reliance on extensive labeled data while addressing the scalability issues in large-scale solar installations. It is based on mean-teacher approach [44] and efficiently leverages unlabeled data. The proposed method enhances model accuracy and generalization by enforcing consistency between the predictions of a student model and a temporally averaged teacher model. Such an approach is crucial for reducing reliance on extensive labeled datasets, addressing a key challenge in the semantic segmentation of PV module defects. To the best of our knowledge there is no current existing work in Photovoltaics making use of semi-supervised learning approaches or reduced labeled data for defect detection.
@@
Figure 1: An overview of the proposed semi-supervised approach for PV defect detection: It utilizes both labeled and unlabelled images in a semi-supervised learning (SSL) framework using supervised and unsupervised consistency losses to train the defect detection model (top row) and during testing the trained model predicts the defect class for each pixel in EL images (bottom row).
@@
    We demonstrate that the proposed approach outperforms existing state-of-the-art fully supervised methods with merely 20% annotations.
@@ 2 Related Work
Semi-supervised Segmentation
Recently, semi-supervised learning (SSL) techniques [44, 41] have demonstrated strong potential in segmentation tasks by leveraging both limited labeled and abundant unlabeled data. Approaches such as pseudo-labeling, consistency regularization, and teacher-student frameworks have been effectively applied in domains like medical imaging and industrial defect inspection [34, 53, 6, 30], enabling improved performance with minimal supervision. Motivated by these successes, we investigate a semi-supervised segmentation framework tailored for photovoltaic (PV) defect detection using electroluminescence (EL) imagery. To the best of our knowledge, this is the first work on SSL-based segmentation in the context of EL-based PV module inspection. By integrating semantic segmentation with a teacher-student framework, our approach enables effective learning from both labeled and unlabeled data, reducing the reliance on extensive annotations while improving defect localization performance.
@@ 3.2.2 SemiCE as Consistency Loss
$\mathcal{L}_{\text{pos}}=-\frac{1}{N_{\text{U}}}\sum_{i=1}^{N_{\text{U}}}\sum_{j=1}^{H\times W}\begin{cases}Y_{\text{L},ij}\cdot\log\left(\frac{\exp(P_{\text{$\theta$},ij})}{\sum_{k=1}^{H\times W}\exp(P_{\text{$\theta$},ik})}\right),&\text{if }Y_{\text{L},ij}=1\text{ and }P_{\text{$\theta$},ij}\geq t_{\text{pos}}\\ 0,&\text{otherwise}.\end{cases}$ (2)
@@
$\mathcal{L}_{\text{neg}}=-\frac{1}{N_{\text{U}}}\sum_{i=1}^{N_{\text{U}}}\sum_{j=1}^{H\times W}\begin{cases}(1-Y_{\text{L},ij})\cdot\log\left(\frac{\exp(-P_{\text{$\theta$},ij})}{\sum_{k=1}^{H\times W}\exp(-P_{\text{$\theta$},ik})}\right),&\text{if }Y_{\text{L},ij}=0\text{ and }P_{\text{$\theta$},ij}<-t_{\text{neg}}\\ 0,&\text{otherwise}.\end{cases}$ (3)
Here, $N_{\text{U}}$ represents the total number of unlabeled images, $P_{\text{$\theta$},ij}$ represent the logits predicted by the model, targets are the ground truth labels, $Y_{\text{L},ij}$ is the ground truth label for the $i$-th sample and $j$-th class, and $H\times W$ represents the spatial dimensions of the image. The final Semi Cross Entropy Loss $\mathcal{L}_{\text{SemiCE}}$ is the sum of $\mathcal{L}_{\text{pos}}$ and $L_{\text{neg}}$: The SemiCE loss is the same as the Consistency loss ($\mathcal{L}_{\text{cons}})$ for the student-teacher framework where this loss component tries to make the predictions of the student model consistent,
$\mathcal{L}_{\text{cons}}=\mathcal{L}_{\text{SemiCE}}=\mathcal{L}_{\text{pos}}+\mathcal{L}_{\text{neg}}.$ (4)
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sup}}+\lambda\mathcal{L}_{\text{cons}},$ (5)
where $\lambda$ is the weighting coefficient that determines the importance of the consistency loss relative to the supervised loss. Section  4.2 further explains on the class imbalance selection of positive ($t_{\text{pos}}$) and negative ($t_{\text{neg}}$) thresholds along with how different values of these threshold affect the results
@@ 4 Experiments and Evaluation
Evaluation Metrics
To evaluate the performance of our proposed method, we use four different metrics: Mean Intersection over Union (mIoU), Precision, Recall, and F1 score. Precision quantifies the proportion of predicted defect pixels that are correctly identified, and is defined as:
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}},$ (6)
where TP (true positives) denotes the number of correctly predicted defect pixels, and FP (false positives) denotes the number of non-defect pixels incorrectly predicted as defects. Recall measures the ability of the model to identify all actual defect pixels, and is given by:
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}},$ (7)
where FN (false negatives) corresponds to defect pixels that were not detected by the model. F1 score, the harmonic mean of precision and recall, offers a balanced metric that considers both false positives and false negatives:
@@
Mean IoU (mIoU) evaluates the overlap between predicted segmentation masks and ground truth annotations, averaged across all classes. It is defined as:
$\text{mIoU}=\frac{1}{N_{\text{classes}}}\sum_{i=1}^{N_{\text{classes}}}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}},$ (9)
where $N_{\text{classes}}$ is the total number of classes, and $\text{TP}_{i}$, $\text{FP}_{i}$, and $\text{FN}_{i}$ represent the true positives, false positives, and false negatives for class $i$, respectively. These metrics provide a comprehensive assessment of segmentation accuracy, particularly in the context of defect detection in EL imagery. Together, these metrics provide a comprehensive assessment for accuracy, completeness, and reliability of the segmentation model in identifying and localizing defects in photovoltaic modules.
@@ 4.1 Quantitative Evaluation
Regarding the performance on UCF-EL dataset with the semi-supervised learning scenario with 20% labeled data, we record average values for metrics as mIoU of 67%, precision of 72.50%, recall of 88.90%, and F1-score of 78.17% whereas with 20% labeled data in the supervised learning yielded mIoU of 8.76%, precision of 25.79%, recall of 26.27% and F1-score of 14.73%. Based on the results we can observe that semi-supervised learning with PV-S3 has yielded significant improvement in results as compared to fully supervised learning with same amount of labeled dataset. The scores with the baseline model of supervised learning with 20% labeled data and semi-supervised learning with 20% labeled data has been summarized in Table  1. To further evaluate the robustness of PV-S3, we perform three experimental runs with different subsets of labeled samples which are randomly selected from the training dataset. The mean and standard deviation of performance metrics for each defect class are shown in Table  2. We observe that the proposed SSL approach significantly outperform the supervised baseline in all metrics consistently, demonstrating its robustness.
mIoU Precision Recall F1-Score
@@
mIoU Precision Recall F1-Score
@@
With 100% labeled data using fully supervised learning, the Deeplabv3+ model achieves mean metrics of mIoU: 70.84%, precision: 80.12%, recall: 86.29%, and F1-score: 82.03% (Table  3). Comparing these results with our 20% labeled semi-supervised learning approach, PV-S3 yields comparable results to fully supervised learning, with even higher precision. Leveraging 100% labeled data in the semi-supervised framework with PV-S3 significantly improves performance, yielding mean mIoU: 74.75%, precision: 79.96%, recall: 91.30%, and F1-score: 84.24%. This improvement underscores the benefits of a larger dataset in better understanding and segmenting defects in PV module EL images.
mIoU Precision Recall F1-Score
@@
Table 3: Comparison of PV-S3 with our fully supervised method on UCF-EL dataset: The table shows the performance of different defect classes under Semi Supervised Learning (SSL) with 100% labeled images, and Fully Supervised Learning (FSL) with 100% labeled images. The metrics evaluated include mIoU, Precision, Recall, and F1-Score, facilitating a direct comparison of the impact of labeling extent on model performance.
Baselines Labeled Data (%) mIoU (%) Precision (%) Recall (%) F1-Score (%)
@@
Comparisons with existing Fully Supervised learning on UCF-EL dataset, performed by [18], across key metrics are summarized in Table  4. We observe improved performance with merely 20% labeled samples and significant enhancement in mIoU, Precision, and F1-Score using all labeled samples. Our approach, utilizing just 20% of labeled samples, yields a notable absolute enhancement, including a 9.7% increase in mIoU, a 13.5% rise in Precision, a 29.15% boost in Recall, and a 20.42% improvement in F1-Score compared to the previous state-of-the-art supervised method, reducing annotation expenses by 80%.
Experiment mIoU (%) Precision (%) Recall (%) F1-Score (%)
@@
On the Combined Solar Benchmark Dataset (CSB-Dataset) we evaluate PV-S3 on all three subsets with three major defect classes Crack, Gridline and Inactive. The results for four key metrics, mIoU, Precision, Recall and F1-Score, are summarized in Table  5. We observe competitive performance with PV-S3 using only 20% labeled samples across all metrics when compared with fully supervised method using 100% labeled samples. We also compare the performance of PV-S3 with existing methods on these datasets. This comparison is shown in Table  6. The performance of PV-S3 is compared with various baseline models which include PSPNet, UNet-12 and UNet-25 using both 20% (Our-SSL) and 100% (Our-FSL) labels. We observe that conventional semantic segmentation models like PSPNet and UNet variants which are UNet-12 and UNet-25 performed poorly in terms of the Intersection of Union as compared to Deeplab variants. In terms of recall score, UNet-25 has yielded the highest recall score of 60.66% while PSPNet yielded the lowest score of 43.0%. We observe that PV-S3 outperforms other baseline models in terms of all metric scores. Using merely 20% of labeled images for training we observed an increase of 30% in IoU with a reduction in 80% annotation cost compared to the state-of-the-art model on CSB-Dataset, further using our model with 100% labeled images an increase of 35.67% in mIoU and 0.33% increase in recall was observed with state-of-the-art model on CSB-Dataset.
Baseline Models Labeled Data (%) mIoU (%) Precision (%) Recall (%) F1-Score (%)
@@ 4.2 Discussion and Analysis
mIoU Precision Recall F1-Score
Defect Class MSE SemiCE-S1 SemiCE-S2 MSE SemiCE-S1 SemiCE-S2 MSE SemiCE-S1 SemiCE-S2 MSE SemiCE-S1 SemiCE-S2
@@
Detecting defects in PV modules presents a significant challenge due to class imbalance: certain defects, such as Interconnect and Corrosion, occur less frequently in the dataset compared to defects like Crack and Contact. To address this imbalance, we employ the Semi Cross-Entropy (SemiCE) loss function, which uses adaptive thresholds for positive and negative predictions during training. The positive threshold ($t_{pos}$) controls how confidently a prediction must be classified as a defect, while the negative threshold ($t_{neg}$) ensures that only strong negative predictions contribute to the loss. This threshold-based filtering helps counter the common issue in imbalanced datasets where the model tends to overpredict frequent classes and underpredict rare ones due to skewed confidence distributions. By lowering $t_{pos}$, the model becomes more receptive to low-confidence predictions, improving recall for minority classes. Conversely, a higher $t_{pos}$ enforces stricter confidence, favoring precision. The negative threshold further ensures that noisy or ambiguous negatives do not dominate the learning signal, helping stabilize training across all classes, including those with fewer examples.
@@
In our experiments, we varied both the SemiCE positive and negative thresholds to observe their impact on Interconnect defect class performance. Performance with different thresholds is shown in Table 8, with original results in Table  1 using initial thresholds of 0.6 (positive) and 0 (negative), chosen for their balanced nature. We conducted experiments by varying the thresholds to observe their impact on model performance across multiple defect types, changing the thresholds significantly affects not only Interconnect defects but also Crack, Contact, and Corrosion classes. For instance, a lower positive threshold ($t_{pos}=0.2$) and moderate negative threshold ($t_{neg}=0.4$) yielded improved performance across all defect classes, with Interconnect achieving an mIoU of 64% and notable improvements in F1-scores for Crack and Contact defects as well. Through experimentation, we aimed to identify threshold values optimizing the metrics while maintaining model consistency, shedding light on the complex relationship between prediction confidence and model performance.
These observations suggest that adjusting thresholds helps the model better capture difficult defects (such as Interconnect) by improving performance, while still maintaining high performance for more frequent classes (such as Crack). Thus, the SemiCE loss function, through careful threshold tuning, effectively mitigates problem of detecting difficult defects across different defect types in the dataset. Based on our experiments, we chose the thresholds ($t_{pos}=0.2$, $t_{neg}=0.4$) to achieve the best overall balance between precision and recall, while maintaining model consistency.
Threshold Crack Contact Interconnect Corrosion Average
Pos, Neg mIoU Prec Rec F1 mIoU Prec Rec F1 mIoU Prec Rec F1 mIoU Prec Rec F1 mIoU Prec Rec F1
0.0, 0.0 89 95 93 94 51 65 71 68 50 71 65 68 92 98 94 96 71 82 81 82
0.2, 0.4 89 96 92 94 53 63 78 70 64 74 83 78 81 91 88 89 72 81 85 83
0.3, 0.2 89 98 91 94 52 61 77 68 44 50 79 61 67 69 96 80 63 70 86 76
0.4, 0.2 91 97 93 95 52 64 73 68 45 48 81 62 76 77 98 86 66 72 86 78
0.5, 0.5 91 98 93 95 53 64 75 69 46 46 97 63 85 87 97 92 69 74 91 80
0.6, 0.0 89 98 90 94 49 57 79 66 40 43 89 58 90 92 97 95 67 73 89 78
@@
We evaluated PV-S3’s performance using different proportions of labeled data (5% to 20%) and compared it with fully supervised learning using 100% labeled data. Results in Figure  8 indicate that semi-supervised learning with 5% labeled data initially underperforms compared to fully supervised learning. However, performance notably improves at 10% labeled data and continues to increase gradually up to 20%, approaching scores achieved with 100% labeled data for mIoU, Recall, and F1 score. Notably, recall scores for semi-supervised learning even exceed those of fully supervised learning. Further we observe that the recall score for the semi-supervised model trained with 20% labeled data and 80% unlabeled data exceeds that of the model trained with 100% labeled data. Usually, semi-supervised learning methods which utilized unlabelled samples outperform fully supervised methods [50, 41, 51], This counterintuitive result can be attributed to the benefits of semi-supervised learning in leveraging unlabeled data to improve model generalization. The inclusion of unlabeled data acts as a regularizer, preventing the model from overfitting to the labeled data, which might contain noise or biases. In contrast, training on 100% labeled data without the additional unlabeled data may lead to overfitting, especially if the labeled data is limited in diversity.By incorporating a large amount of unlabeled data, the semi-supervised model learns more robust feature representations, capturing the underlying structure of the data more effectively. The semi-supervised approach allows the model to generalize better to unseen data, resulting in higher recall by correctly identifying more true positive instances of defects.
@@
We analyze the training process of PV-S3 with epochs and compare with the supervised approach. The plots for different metrics are shown in Figure  9. Here we observe that for supervised learning with only 20% images (SL (20%)), due to less number of images and class imbalance, the model is not able to learn the features accurately and hence it has very low pixel accuracy and high loss values and have low values of precision and recall as well. For PV-S3, we observe an increasing trend for pixel accuracy, precision, and recall and decreasing trend for loss as training progresses.
Figure 9: Analyzing model training: The top row plots show the variation of pixel accuracy (left), loss (middle), and mIoU (right) with varying training epochs, and the bottom row shows the performance metrics variation with epochs for supervised and semi-supervised learning with 20% images for precision (left), recall (middle), and F1-score (right).